# DCFormer: Efficient 3D Vision-Language Modeling with Decomposed Convolution

### v1

@@ Methods
Processing 3D CT images presents significant computational challenges due to their high resolution and volumetric complexity. 3D convolution is computationally expensive due to its cubic scaling with kernel size, and the self-attention mechanism further amplifies the burden by scaling quadratically with input size. To address these challenges, we propose DCFormer, a hybrid architecture that decomposes 3D convolution into multiple 1D components, significantly reducing computational overhead. DCFormer adopts MetaFormer [33, 36] as its backbone and employs decomposed convolutions as the token mixer for feature extraction. Its hierarchical structure further enhances feature representation in 3D CT images. Finally, the DCFormer-based image encoder is integrated into a CLIP framework for image-text alignment. As demonstrated in our results, this decomposition strategy substantially reduces parameter count and computational cost while maintaining high performance in both zero-shot and fine-tuning settings.
Formulation of DCFormer block
The core concept, MetaFormer [33, 36], is a general architecture abstracted from transformers [37]. In the Metaformer architecture, the token mixer (e.g. self-attention, depthwise convolution) is not specified, while normalization, a channel mixer (e.g. MLP), and residual connections [38] are retained. The input $I$ is first passed into a patch embedding block, such as a convolution:
@@
An illustration of the DCFormer and its Pytorch-like implementation are shown in Figure 3 and Algorithm 1, respectively.
@@
Given an input tensor $X\in\mathbb{R}^{1\times C\times H\times W\times D}$ with batch size $B=1$ a 3D depthwise convolution [29] with a kernel size ($k_{h},k_{w},k_{d}$) where $k_{h}=k_{w}=k_{d}=k$, the total parameter count and FLOPs can be calculated as $Ck^{3}$ and $2CHWDk^{3}$, respectively (bias is omitted for simplicity). Note that both parameters and FLOPS increase cubically as $k$ increases. This introduces extensive computational burden, especially for large kernel sizes and when these convolutions are used repeatedly in an end-to-end network, where the effect of the cubic scaling compounds over multiple layers.
By decomposing the 3D depthwise convolution into three 1D depthwise convolutions, the parameter count and FLOPS are significantly reduced to $3Ck$ and $6CHWDk$, respectively, as each 1D convolution operates independently along a single spatial dimension rather than all three simultaneously. A comparison of FLOPs and parameter counts for 3D depthwise and decomposed convolutions is shown in Figure 2. For the decomposed convolution, the computational cost scales linearly with respect to kernel size $k$, as opposed to the cubic scaling in the 3D depthwise convolution.
When the kernel size is larger than 3, the complexity increase becomes much more significant. Large kernel sizes such as 7, 11 [31, 32] and even up to 51 [43, 44] have been shown to improve model performance. In 3D imaging, our decomposition approach becomes crucial, particularly in deep networks where multiple convolution layers are applied.
Figure 4: Overview of DCFormer
@@
Based on the above findings, we formulate DCFormer as the token mixer and develop a family of architectures of varying sizes, using DCFormer as the main building block. Figure 4 illustrates the general architecture of our proposed DCFormer. Following previous work [38, 45, 31, 33, 46, 32], DCFormer adopts a hierarchical structure with four stages and an initial stem stage following [46, 47].
In the stem stage, we use a decomposed convolution with a kernel size of 7 and a stride of 4. The downsampled images are then processed with three additional decomposed convolutions, each with a kernel size of 3 and a stride of 1. The outputs of the stem stage are passed into the hierarchical structure, where each stage produces tokens of sizes $\frac{H}{8}\times\frac{W}{8}\times\frac{D}{8}$, $\frac{H}{16}\times\frac{W}{16}\times\frac{D}{16}$, $\frac{H}{32}\times\frac{W}{32}\times\frac{D}{32}$, and $\frac{H}{64}\times\frac{W}{64}\times\frac{D}{64}$, respectively, where $H$, $W$, $D$ represent height, width and depth of the input image volume.
We propose four model variants: nano, naive, tiny, small. The number of layers in each stage are [1, 1, 1, 1] for nano, [2, 2, 2, 2] for naive, [2, 3, 3, 2] for tiny and [2, 3, 6, 2] for small models. Additionally, each stage begins with patch embedding where a 3D max pooling operation with a kernel size of 3 and a stride of 2 is used to reduce the image dimensions [47]. Following [7, 31, 32], we set the MLP ratio to 4, while the kernel size for the decomposed convolution is set to 7, as in [45, 31]. Further configuration details are provided in Table 2.
@@
Nano Naive Tiny Small
@@ 4 H/64 X W/64 X D/64 Patch Embedding Patch size 3 x 3 x 3, stride 2
Parameters (M) 0.92 5.85 15.1 26.86
FLOPS(G) 34.21 49.48 168.2 202.9
Table 2: DCFormer configuration details.
@@
Figure 5: CLIP a) training and b) zero-shot inference framework
Zero-shot multi-abnormality detection with DCFormer
For zero-shot multi-abnormality detection, we follow the same approach as CheXzero[21] and CT-CLIP[26]. Specifically, two similarity scores are computed using the cosine similarity between the embedding of the CT image of interest and the embeddings of a positive (’{Pathology} is present.’) prompt and a negative (’{Pathology} is not present.’) prompt. A softmax function is then applied to these similarity scores to estimate the likelihood of each abnormality being present in the given CT image (see Figure. 5 (b) for illustration). It should be noted that the zero-shot performance can potentially be improved by experimenting with different prompt variations. For instance, possible prompt templates including: ’Signs of {Pathology} are/are not detected in the image.’, ’There is/is not {Pathology}.’, ’{Pathology} is/is not observed.’ ’Indication/No indication of {Pathology}.’. Although prompt engineering plays a critical role in improving image-text alignment, we adopt the prompt structure proposed by Hamamci et al. [26] (’{ Pathology} is/is not present.’) to maintain consistency with prior work and ensure a fair comparison.
@@
For model training, we follow the same data splitting and pre-processing strategy as in [26]. Specifically, we use 20,000 patients for training and 1,304 for validation. For pre-processing, we first resize each CT volume to a spacing of 0.75 mm on the x-axis and y-axis and 1.5 mm on the z-axis. Then, the volumes are center-cropped or padded to a fixed size of 512 x 512 x 256. Finally, we clip the Hounsfield Unit (HU) values of each CT image volume to [-1000,1000] and normalize them to [-1,1]. We train models using the AdamW optimizer [48] with an initial learning rate $1^{-5}$. We dot not apply learning rate scheduling or warmup, as we have not observed any significant improvements. All models are trained for 15 epochs.
@@ Acknowledgements
This work was supported by the Department of Medicine and the Intelligent Clinical Care Center at the University of Florida College of Medicine.
Author contributions statement
G.C.A. conducted study design, software development, experimental setup, data analysis, visualization, and manuscript writing. K.G. provided numerous insights and contributed to discussions throughout this study. W.S. oversaw the project, conducted study design, experimental setup, and data analysis, and contributed to manuscript writing. All authors reviewed the manuscript.
@@
Competing interests The authors have no conflict of interest to declare.

### v2

@@ Methods
Processing 3D images presents significant computational challenges due to their high resolution and volumetric complexity. The computational complexity of standard 3D convolutions scales cubically with kernel size, and self-attention mechanisms further amplify the burden by scaling quadratically with image size. To address these challenges, we developed DCFormer, a hybrid architecture that decomposes 3D convolutions into multiple 1D components, significantly reducing computational overhead. DCFormer adopts the MetaFormer paradigm [32, 37] as its backbone and employs decomposed convolutions as the token mixer for feature extraction. Its hierarchical design (described below) further enhances feature representation in 3D images. Finally, the DCFormer-based image encoder is integrated into a CLIP framework for image-text alignment. As demonstrated in our results, this decomposition strategy substantially reduces parameter count and computational cost while maintaining high performance.
Formulation of the DCFormer block
Our design builds on the MetaFormer architecture [32, 37], which is a general vision transformer paradigm. In the Metaformer framework, the token mixer (e.g. self-attention, depthwise convolution) is not specified, while normalization, a channel mixer (e.g. MLP), and residual connections [38] are retained. The input image tensor $I$ is first passed through a patch embedding block, such as a convolution:
@@
An illustration of the DCFormer block and its relationship to existing architectures is shown in Figure 3, and PyTorch-style pseudocode for the Decomposed Convolution operation is provided in Algorithm 1, respectively.
@@
Given an input tensor $X\in\mathbb{R}^{C\times H\times W\times D}$, a standard 3D depthwise convolution [29] with a kernel size $k$ in all dimensions has $Ck^{3}$ parameters and $2CHWDk^{3}$ FLOPs (bias operations are omitted for simplicity). This cubic growth in complexity can quickly become prohibitive for large $k$ and for deep networks, where multiple such layers compound the cost. In contrast, our decomposed 3D convolution uses three 1D depthwise convolutions, which collectively require only $3Ck$ parameters and $6CHWDk$ FLOPs. Thus, the parameter count and computational cost scale linearly with kernel size rather than cubically. Figure 2 compares the theoretical FLOPs and parameter counts of a standard 3D depthwise convolution with those of a decomposed convolution, highlighting the drastic efficiency gain of the latter. Notably, for any kernel size larger than 3, the difference in complexity becomes increasingly significant. Large kernel sizes (e.g., 7 or 11, as used in prior works [31, 43], and even 51 as explored in recent studies [44, 45]) have been shown to improve CNN performance. In 3D imaging, our decomposition approach makes the use of such large kernels feasible, which is particularly important in deep networks where multiple convolutional layers are stacked.
Figure 4: Hierarchical architecture of DCFormer.
@@
Using the DCFormer block as the core component, we constructed a hierarchical vision encoder architecture (Figure 4). Following prior work [38, 46, 31, 32, 47, 43], DCFormer includes an initial stem stage followed by four hierarchical stages. In the stem stage, we use a decomposed convolution with a kernel size of 7 and a stride of 4. The downsampled images are then processed with three additional decomposed convolutions, each with a kernel size of 3 and a stride of 1. The output of the stem is then fed into four stages of DCFormer blocks, where each stage produces tokens of sizes $\frac{H}{8}\times\frac{W}{8}\times\frac{D}{8}$, $\frac{H}{16}\times\frac{W}{16}\times\frac{D}{16}$, $\frac{H}{32}\times\frac{W}{32}\times\frac{D}{32}$, and $\frac{H}{64}\times\frac{W}{64}\times\frac{D}{64}$, respectively, where $H$, $W$, $D$ represent height, width and depth of the original image volume.
We implemented DCFormer in three model variants: nano, naive, and tiny. The number of layers in each stage is [1, 1, 1, 1] for nano, [2, 2, 2, 2] for naive, and [2, 3, 3, 2] for tiny. At the beginning of each stage (except the stem), we include a patch embedding downsampling operation implemented via 3D max pooling with a kernel size of 3 and a stride of 2, to further reduce the feature map size [48]. We set the MLP expansion ratio to 4 in all DCFormer blocks, following [7, 31, 43], and use a base decomposed convolution kernel size of 7, consistent with the design in [46, 31]. Additional configuration details for each model variant are provided in Table 4.
Table 4: Architectural configurations of the DCFormer model variants.
@@
Nano Naive Tiny
@@ 4 H/64 X W/64 X D/64 Patch Embedding Patch size 3 x 3 x 3, stride 2
Parameters (M) 0.92 5.85 15.1
FLOPS(G) 34.21 49.48 168.2
@@
Figure 5: DCFormer-based CLIP framework: (a) Training with paired CT volumes and reports, (b) Zero-shot inference with text prompts, (c) Fine-tuning for multi-label classification, and (d) Text-to-image retrieval based on embedding similarity.
@@
For model training, we followed the same data splitting and pre-processing strategy as in [26]. Specifically, we used 20,000 patients for training and 1,304 for validation. For pre-processing, we first resized each CT volume to a spacing of 0.75 mm on the x- and y-axes, and 1.5 mm on the z-axis. The volumes were then center-cropped or padded to a fixed size of 512 × 512 × 256. Finally, we clipped the Hounsfield Unit (HU) values of each CT image volume to the range $[-1000,1000]$ and normalized them to $[-1,1]$. We trained our models using eight NVIDIA A100 80GB GPUs. For optimization, we used AdamW [49] with an initial learning rate of $10^{-5}$. We did not apply learning rate scheduling or warmup, as we did not observe any significant improvement. All models were trained for 15 epochs. For fine-tuning, we integrated a linear classification layer into the image encoder and trained only this layer while keeping the image encoder frozen.
@@ Acknowledgements
This work was supported by the Department of Medicine and the Intelligent Clinical Care Center at the University of Florida College of Medicine. The authors express their sincere gratitude to the NVIDIA AI Technology Center at the University of Florida for their invaluable feedback, technical guidance, and support throughout this project.
Author Contributions
G.C.A. conducted the study design, software development, experimental setup, data analysis, visualization, and manuscript writing. Y.X. contributed to the experimental setup, manuscript writing, and discussions. K.G. provided valuable insights and contributed to discussions throughout the study. W.S. supervised and conceptualized the project, guided methodological development and experimental setup, and contributed to writing the original manuscript. All authors reviewed and approved the final manuscript.
@@
Competing interests: The authors have no conflict of interest to declare.